Adv. Math. Econ. 8, 433-447 (2006)

Licensing agreements as bargaining outcomes: general results and two examples

Naoki Watanabe^{1} and Shigeo Muto^{2}

1 Guraduate School of Economics, Hitotsubashi University, 2-1 Naka, Kunitachi, Tokyo 186-8601, Japan

2 Department of Social Engineering, Graduate School of Decision Science and Technology, Tokyo Institute of Technology, 2-12-1 Ookayama, Meguro, Tokyo 152-8552, Japan

Received: September 7, 2005 Revised: September 29, 2005

JEL classification: D45, D43, C71

Abstract. We investigate coalition structures formed by an external licensor of a patented innovation and firms operating in oligopolistic markets, and study licensing agreements reached as the bargaining outcomes under those coalition structures. The following are shown. (I) If the goods are not so substitutive in Bertrand competition, the grand coalition forms and the core is not empty, whereas in Cournot competition the core is empty under any coalition structure. (II) When the optimal number of licensees is one in the Bertrand duopoly with differentiated goods, the bargaining set gives more revenue to the licensor than licensing by means of upfront fee expects to give.

Key words: licensing, oligopoly, stable coalition structure, core, bargaining set